# Adaptively Robust and Sparse

### v1

@@ Introduction
Table 6 summarizes the result of the real-life dataset comparison. The reported values are CERs for each approach. In order to better illustrate the performance of that method, we have emphasized the first and second best CER values in bold when applied to a certain real-life dataset. While traditional methods such as KC, PCA-KC, and TKM prove effective in certain contexts, they generally yield higher CERs across the datasets compared to the proposed ARSK algorithm and its variants. For instance, in the glass and Breast Cancer dataset, the CER values obtained by KC, PCA-KC, and TKM are considerably higher than those of the best-performing ARSK. In seven real-world datasets, our ARSK method ranked among the top two best methods in six of the datasets. Therefore, the proposed ARSK algorithm consistently achieves superior performance with lower CERs across most of the real datasets, particularly when using soft-thresholding to obtain the error matrix $E$. ARSK demonstrates competitive performance in this regard.
@@
This comparative analysis emphasizes the ARSK algorithm’s success in significantly reducing classification error rates across diverse real-world datasets, showcasing the versatility and adaptability of our proposed clustering method. Through its different configurations (e.g., soft-soft, soft-SCAD, SCAD-soft and SCAD-SCAD), allow it to handle various data structures and sizes effectively. The consistent performance across different real-world datasets underscores the broad applicability of our approach in various fields.
@@
In this paper, we propose an Adaptively Robust and Sparse $K$-means Clustering algorithm designed to handle multiple tasks within a unified framework for a comprehensive analysis of contaminated high-dimensional datasets. Our approach exhibits greater flexibility in handling outliers within real-world datasets. We also proposed a modified Gap statistic with an alternating optimization algorithm to search for tuning parameters. This approach significantly reduces the computational time required for parameter tuning compared to traditional methods. The experiment results demonstrate the significant capabilities of the proposed approach in revealing group structures in high-dimensional data. Our model not only performs well in detecting outliers but also identifies significant variables directly, leading to a more thorough understanding of complex datasets.
In our current study, we assume that the number of clusters is known, consistent with the traditional $K$-means method. Additionally, Chen and Witten (2023) proposed a test for detecting differences in means between each cluster estimated from $K$-means clustering. This technique can be applied to the modified dataset $x_{ij}-E_{ij}$, where noise variables (i.e., $w_{j}=0$) are removed, to help determine which clusters should be retained.
Furthermore, We have consistently utilized a single penalty factor $\lambda_{1}$ in the objective function (4). This ensures a uniform approach to detecting outliers across various groups. However, given the complexity and diversity of the data, a more sophisticated strategy is needed. This might involve using different penalty factors, or distinct group penalty functions for different groups. One might employ $\sum_{k}\sum_{i}P_{k}(E_{i\cdot};\lambda_{k})$ in (4). Incorporating this approach into unsupervised learning presents challenges in selecting not just the penalty factor but also the group penalty function. These extensions are left to our future work.
Acknowledgement
@@
Appendix
Proofs solution to (9)
@@
$L(w_{j},\lambda_{2})=\sum_{j=1}^{p}W_{j}Q_{j}-\sum_{j=1}^{p}(P_{2}(w_{j};\lambda_{2})+\frac{1}{2}w_{j}^{2}),$
where $Q_{j}=Q_{j}^{R}(x_{\cdot j};\Theta,E_{\cdot j})$.
Taking the derivative with respect to $w_{j}$ and setting it to zero, we obtain:
$Q_{j}-w_{j}-\Gamma_{j}=0$
When the $L_{1}$ penalty $P_{2}(w_{j};\lambda_{2})$ is applied, $\Gamma_{j}$ represents the subgradient of the $L_{1}$ penalty with respect to $w_{j}$. Solving the above equation yields:
@@
When the SCAD penalty $P_{2}(w_{j};\lambda_{2})$ is applied, $\Gamma_{j}$ represents the subgradient of the SCAD penalty with respect to $w_{j}$. Solving the above equation yields:
$S^{\text{SCAD}}(x;\lambda_{2})=w_{j}=\left\{\begin{array}[]{ll}\text{sgn}(Q_{j})(|Q_{j}|-\lambda_{2})_{+}&\text{if}|Q_{j}|\leq 2\lambda_{2}\\ \frac{(a-1)Q_{j}-a\lambda_{2}\cdot\text{sign}(Q_{j})}{(a-1)-1}&\text{if}2\lambda_{2}\leq|Q_{j}|<a\lambda_{2}\\ Q_{j}&\text{if}a\lambda_{2}\leq|Q_{j}|\end{array}\right.$
Project $\hat{w_{j}}$ onto the parameter space $\sqrt{\sum_{j=1}^{p}w_{j}}=1$. The final solution can be rewritten as:
$w_{j}=\frac{S(Q_{j};\lambda_{2})}{\sqrt{\sum_{j=1}^{p}(S(Q_{j};\lambda_{2}))^{2}}}$

### v2

@@ 3.4.1 UCI data application
Table 6 summarizes the result of the real-life dataset comparison. The reported values are CERs for each approach. In order to better illustrate the performance of that method, we have emphasized the first and second best CER values in bold when applied to a certain real-life dataset. While traditional methods such as KC, PCA-KC, and TKM demonstrate effectiveness in certain contexts, they generally yield higher CERs across the datasets compared to the proposed ARSK algorithm and its variants. For instance, in the glass and Breast Cancer dataset, the CER values obtained by KC, PCA-KC, and TKM are considerably higher than those of the best-performing ARSK. In seven real-world datasets, our ARSK method ranked among the top two best methods in six of the datasets. Therefore, the proposed ARSK algorithm consistently achieves superior performance with lower CERs across most of the real datasets, particularly when using the multivariate soft-threshold operator to obtain the error matrix $\bm{E}$. ARSK demonstrates competitive performance in this regard. Furthermore, Figure 1 summarizes the number of zero elements in the variable weight vector of each method, where ARKC adopted a combination of thresholding corresponding to the minimum CER. Figure 1 shows that when analyzing ultra-high-dimensional data, the variable weight vector predicted by ARSK exhibits strong sparsity.
Figure 1: The number of zero elements in the variable weight vector of each method, with results ranging from low-dimension to high-dimension dataset
@@
This comparative analysis emphasizes the ARSK algorithm’s success in significantly reducing classification error rates across diverse real-world datasets, showcasing the versatility and adaptability of our proposed clustering method. Its different configurations (e.g., soft-soft, soft-SCAD, SCAD-soft, and SCAD-SCAD) allow it to handle various data structures and sizes effectively. The consistent performance across different real-world datasets underscores the broad applicability of our approach in various fields.
3.4.2 Analyzing the group structure of recognition of handwritten digits
The dataset was downloaded from Kaggle, a data science competition platform. It consists of digitized images of the numbers 0 to 9 handwritten by 13 subjects. The images were divided into 64 non-overlapping blocks of $4\times 4$ bits. For each block, we observe the number of black pixels. Hence, each image is represented by $p=64$ variables recording the count of pixels in each block, taking integer values between 0 and 16. Its row names identify the true digit in the image, and the column names identify the position of each block in the original bitmap.
Figure 2: The images of handwritten digits data
Figure 3: The heatmap of the variable weight vector obtained by SCAD-soft ARSK
Our focus is to detect an entire variable weight and group structure, i.e., ten groups. Given the superior performance of SCAD-soft ARSK in subsection 3.4.1, we apply RSKC ($\alpha=0.1$), RSKC ($\alpha=0.2$), WRCK, and SCAD-soft ARSK to this dataset. Although the precise contribution of individual pixels to the clustering process remains indeterminate, visual inspection of the real images of handwritten digits data (0 and 9) as Figure 2 shows that the pixels situated in both the initial and terminal columns exhibit uniform white colouration. Thus, these 16 pixels can be reasonably deemed redundant and inconsequential for the clustering. The optimal robustness and sparsity parameter for ARSK is also selected using the Gap statistic. The evaluation of the resulting clustering solution is presented in Table 7. We also examine the final variable weights obtained by the sparse $k$-means-based algorithms. Table 8 illustrates each method’s number of zero elements within the weight vector derived. The analysis of the weight values reveals that WRCK is ineffective in achieving sparsity within the variable weight vector, as evidenced by the exclusive presence of nonzero elements throughout. Correspondingly, it also exhibits the highest CER. As shown in Table 7, the performance of RSKC and ARSK is similar, with CER values around 0.8. However, the RSKC ($\alpha=0.2$) produces weight vectors with only 6 zero elements, indicating a moderate sparsity level. In contrast, the proposed ARSK approach generates weight vectors with 13 zero elements that are close to 16. We also presented a heatmap, in Figure 3, visualization of the variable weight vector derived from the SCAD-soft ARSK algorithm as Figure 2, wherein the intensity of the chromatic scale correlates positively with the magnitude of the weights. We can conclude that these findings correspond with those shown in Figure 2. These results demonstrate that ARSK exhibits both sparsity and accuracy of feature selection capabilities.
RSKC-0.2 RSKC-0.1 SCAD-soft ARSK WRCK
CER 0.086 0.077 0.079 0.156
Table 7: Evaluation of the clustering performance on recognition of handwritten digits
RSKC-0.2 RSKC-0.1 SCAD-soft ARSK WRCK
CER 6 9 13 0
Table 8: The number of zero elements within the weight vector derived from each method
@@ 3.4.2 Analyzing the group structure of recognition of handwritten digits
In this paper, we propose an Adaptively Robust and Sparse $K$-means Clustering algorithm designed to handle multiple tasks within a unified framework for a comprehensive analysis of contaminated high-dimensional datasets. Our approach demonstrates greater flexibility in handling outliers within real-world datasets. We also proposed a modified Gap statistic with an alternating optimization algorithm to search for tuning parameters. This approach significantly reduces the computational time required for parameter tuning compared to traditional methods. The experimental results demonstrate the significant capabilities of the proposed approach in revealing group structures in high-dimensional data. Our model not only performs well in detecting outliers but also identifies significant variables directly, leading to a more thorough understanding of complex clustering datasets.
In our current study, we assume that the number of clusters is known and consistent with the traditional $K$-means method. Additionally, Chen and Witten (2023) proposed a test for detecting differences in means between each cluster estimated from $K$-means clustering. This technique can be applied to the modified dataset $X_{ij}-E_{ij}$, where noise variables (i.e., $w_{j}=0$) are removed to help determine which clusters should be retained.
Furthermore, we have consistently utilized a single penalty for outliers in the objective function (4). This implements a uniform approach to detecting outliers across various groups. However, given the complexity and diversity of the data, a more sophisticated strategy may be required. This might involve using distinct group penalty functions for different groups. Incorporating this approach into unsupervised learning presents challenges in selecting group penalty functions. These extensions are left to our future work.
Acknowledgments
@@
Appendix A Appendix: Derivation of the solution to (9)
@@
$L(w_{j},\lambda_{2})=\sum_{j=1}^{p}w_{j}Q_{j}-\sum_{j=1}^{p}\left(P_{2}(w_{j};\lambda_{2})+\frac{1}{2}w_{j}^{2}\right),$
where $Q_{j}=Q_{j}^{R}(\bm{X}_{:,j};\Theta,\bm{E}_{:,j})$. Taking the derivative with respect to $w_{j}$ and setting it to zero, we obtain:
$Q_{j}-w_{j}-\Gamma_{j}=0.$
When the $L_{1}$ penalty $P_{2}(w_{j};\lambda_{2})$ is applied, $\Gamma_{j}$ represents the subgradient of the $L_{1}$ penalty with respect to $w_{j}$. Solving the above equation yields
@@
When the SCAD penalty $P_{2}(w_{j};\lambda_{2})$ is applied, $\Gamma_{j}$ represents the subgradient of the SCAD penalty with respect to $w_{j}$. Solving the above equation yields
$S^{\text{SCAD}}(x;\lambda_{2})=w_{j}=\left\{\begin{array}[]{ll}\text{sgn}(Q_{j})(|Q_{j}|-\lambda_{2})_{+}&\text{if}\ |Q_{j}|\leq 2\lambda_{2}\\ (a-2)^{-1}\big{\{}(a-1)Q_{j}-a\lambda_{2}\text{sign}(Q_{j})\big{\}}&\text{if}\ 2\lambda_{2}\leq|Q_{j}|<a\lambda_{2}\\ Q_{j}&\text{if}\ a\lambda_{2}\leq|Q_{j}|.\end{array}\right.$
Projecting $\hat{w_{j}}$ onto the parameter space $\sqrt{\sum_{j=1}^{p}w_{j}}=1$, the final solution can be rewritten as
$w_{j}=\frac{S(Q_{j};\lambda_{2})}{\sqrt{\sum_{j=1}^{p}(S(Q_{j};\lambda_{2}))^{2}}}.$